Laurel Enterprises expects eamings next year of $3.73 per share and has a 30% retention rate, which it plans to keep constant. Its equity cost of capital is 11%, which is also its expected return on new investment. If ts earnings are expected to grow forever at a rate of 3% per year, what do you estimate the firms current stock price to be? (Hint its next dividend is due in one year.) The current stock price will be$. (Round to the nearest cent.)
Answer #1

Dividend payout ratio=1-retention ratio

=1-0.3=0.7

Hence expected dividend=Earnings next year*Dividend payout ratio

=(0.7*3.73)=$2.611

Current price=D1/(Cost of capital-Growth rate)

=$2.611/(0.11-0.03)

which is equal to

=$32.64(Approx).